1-sin^4 theta/ 1+sin^2 theta= cos^2 theta
Answer by stanbon(75887)

1-sin^4 theta/ 1+sin^2 theta= cos^2 theta
------
Factor:
(1-sin^2(t))(1+sin^2(t))/(1+sin^2(t)) = 1-sin^2(t)
---------
Reduce the fraction on the left side to get:
1-sin^2(t) = 1-sin^2(t)
=========================
Cheers,
Stan H.
-------------------------